The perimeter of a rectangle is 82 yards. The width is 15 yardsless than the length. Find the length and width of the rectangle.

















Answer by DJ_Sammie(24)

l=length
l-15=width
2l+2w=P
p=perimeter
2l+2(l-15)=82
2l+2l-30=82
4l-30=82
+30 +30
4l=112
4l/4=112/4
l=28
length=28yards
width=l-15
width=28-15
width=13yards
The width is 13 yards and the length is 28 yards.